# Claas Voelcker

### v1

@@ Appendix A Proofs
The first proposition relies on the existence of a deterministic mapping, which we prove here as a lemma.
The proof relies heavily on several propositions from Bertsekas and Shreve [1978], which are restated in subsection A.2 for reader’s convenience. Other topological statements are standard and can be find in textbooks such as Munkres [2018].
@@
As $V$ is continuous and $\mathcal{Z}_{\mathcal{X}}$ is compact, it takes a maximum and minimum over the set $\mathcal{Z}_{\mathcal{X}}$. This follows from the compactness of $\mathcal{Z}_{\mathcal{X}}$ and the extreme value theorem. Furthermore $V_{\min}\leq\theta_{V}(z)\leq V_{\max}$ for every $z\in\mathcal{Z}_{\mathcal{X}}$. By the intermediate value theorem over compact, connected spaces, and the continuity of $V$, for every value $V_{\min}\leq v\leq V_{\max}$, there exists a $z\in\mathcal{Z}_{\mathcal{X}}$ so that $V(z)=v$.
Let $h:\mathcal{Z}_{\mathcal{X}}\times\mathcal{Z}_{\mathcal{X}}\rightarrow\mathbb{R}$ be the function $h(z,z^{\prime})=\left|\theta_{V,\mathcal{Z}}(z)-V(z^{\prime})\right|^{2}$. As $h$ is a composition of continuous functions, it is itself continuous. Let $h^{*}(z)=\min_{z^{\prime}\in\mathcal{Z}_{\mathcal{X}}}h(z,z^{\prime})$. By the intermediate value argument presented above, we know that there exist choices for $z^{\prime}$ so that $h^{*}(z)=0$ for all $z\in\mathcal{Z}_{\mathcal{X}}$.
Since $\mathcal{Z}_{\mathcal{X}}$ is compact, $h$ is defined over a compact subset of $\mathcal{Z}$. By Proposition 4, there now exists a measurable function $f^{*}(z)=z^{\prime}$ so that $\min_{z^{\prime}}h(z,z^{\prime})=h(z,f^{*}(z))=0$. Therefore, the function $f^{*}$ has the property that $V(f^{*}(z))=\mathbb{E}_{p}\left[V(z^{\prime})|z\right]$, as this minimizes the function $h$.
@@ A.1 Main propositions
Furthermore, $f^{*}$ is a minimizer of the population IterVAML loss:
@@
The rest of the proof follows standard approaches in textbooks such as Györfi et al. [2002]. First, expand the equation to obtain:
@@
Expanding the square, apply the tower property to the inner term to obtain:
@@
$=$ $2\mathbb{E}\left[\left[V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]\right]\underbrace{\mathbb{E}\left[\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]-V_{\mathcal{X}}(x^{\prime})\middle|x^{\prime}\right]}_{=0}\right].$
Since the statement we are proving only applies to the minimum of the IterVAML loss, we will work with the $\operatorname*{arg\,min}$ of the loss function above. The remaining equation can then be decomposed into a term dependent on $f$ and independent of $f$:
$\operatorname*{arg\,min}_{f}~{}$ $\mathbb{E}\left[\left[V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]\right]^{2}\right]+\mathbb{E}\left[\left[\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]-V_{\mathcal{X}}(x^{\prime})\right]^{2}\right]$
[…]
Finally, it is easy to notice that $V\left(f^{*}\left(\phi(x)\right)\right)=\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]$ by the definition of $f^{*}$ which concludes the proof.
@@
For clarity of presentation denote samples from the real environment as $x^{(n)}$ for the $n$-th sample after a starting point $x^{(0)}$. This means that $x^{(n+1)}$ is drawn from $p\left(\cdot\middle|x^{(n)}\right)$. Similarly, $\hat{x}^{(n)}$ is the $n$-th sample drawn from the model, with $\hat{x}^{(0)}=x^{(0)}$. We will also assume that $\hat{V}=V_{\text{target}}=V^{\prime}$.
All expectations are taken over $x^{(0)}_{i}\sim\mu$ where $\mu$ is the data distribution, $\hat{x}^{(1)}_{i}\sim\hat{p}\left(\cdot\middle|x^{(0)}_{i}\right)$, $x^{(1)}_{i}\sim p\left(\cdot\middle|x^{(0)}_{i}\right)$, and $x^{(2)}_{i}\sim p\left(\cdot\middle|x^{(1)}_{i}\right)$. We use the tower property several times, with the shorthand notation $\mathbb{E}\left[f\left(x^{(0)}_{i},x^{(1)}_{i},x^{(2)}_{i},\hat{x}^{(1)}_{i}\right)\right]=\mathbb{E}\left[\mathbb{E}\left[f\left(x^{(0)}_{i},x^{(1)}_{i},x^{(2)}_{i},\hat{x}^{(1)}_{i}\right)\right]\middle|x^{(0)}_{i}\right]=\mathbb{E}_{x^{(0)}_{i}\sim\mu}\left[\mathbb{E}_{\hat{x}^{(1)}_{i},x^{(1)}_{i},x^{(2)}_{i}\sim p,\hat{p}}\left[f\left(x^{(0)}_{i},x^{(1)}_{i},x^{(2)}_{i},\hat{x}^{(1)}_{i}\right)\right]\middle|x^{(0)}_{i}\right]$
Case 1:
Proof that $V\notin\operatorname*{arg\,min}_{\hat{V}\in\mathcal{V}}\mathbb{E}_{\mathcal{D}}\left[\hat{\mathcal{L}}_{\text{MuZero}}(p,\hat{V};\mathcal{D})\right]$:
Expand the MuZero loss and take its expectation:
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(\hat{p},\hat{V};\mathcal{D})\right]$
@@
$~{}2\mathbb{E}\left[\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)-\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)\right)\left(\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right)\right]+$ (7)
@@
The first term (Equation 6) is the regular squared Bellman residual $\text{BR}^{2}$. The third term (Equation 8) is independent of $\hat{V}$, so we can drop it when analyzing the minimizer.
@@
as the remainder is independent of $\hat{V}$ again. When we use $\hat{V}=V^{\prime}$, apply the tower property, and use the conditional independence of $\hat{x}^{(i)}_{i}$ and $x^{(i)}_{i}$ given $x^{(0)}_{i}$ in the third equality, we obtain
@@
$=$ $~{}\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)-\left[r\left(x^{(1)}_{i}\right)+\gamma\hat{V}\left(x^{(2)}_{i}\right)\right]\right)\right]$
$=$ $~{}\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)^{2}\right]-\mathbb{E}\left[\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)\left[r\left(x^{(1)}_{i}\right)+\gamma\hat{V}\left(x^{(2)}_{i}\right)\right]\right)\right]$
$=$ $~{}\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)^{2}\right]-\mathbb{E}\left[\mathbb{E}\left[\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)\middle|x^{(0)}_{i}\right)\mathbb{E}\left[\left[r\left(x^{(1)}_{i}\right)+\gamma\hat{V}\left(x^{(2)}_{i}\right)\right]\middle|x^{(0)}_{i}\right]\right]\right]$
$=$ $~{}\mathbb{E}\left[\text{Var}\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)\middle|x^{(0)}_{i}\right)\right].$
@@
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(p,\hat{V};\mathcal{D})\right]$
$=$ $~{}\mathbb{E}\left[\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)-\left(\mathcal{T}V\right)\left(\hat{x}^{(1)}_{i}\right)\right)^{2}\right]+\mathbb{E}\left[\text{Var}\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)\middle|x^{(0)}_{i}\right)\right].$
The true value function minimizes the first term, but not the second, therefore it is not the minimizer of the full loss.
Case 2:
Proof that $p\notin\operatorname*{arg\,min}_{\hat{p}\in\mathcal{P}}\mathbb{E}_{\mathcal{D}}\left[\hat{\mathcal{L}}_{\text{MuZero}}(\hat{p},V;\mathcal{D})\right]$.
@@
We will chose a slightly different expansion than before, using $V\left(x^{(1)}_{i}\right)$.
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(\hat{p},V;\mathcal{D})\right]$
$=$ $~{}\mathbb{E}\left[\left[V\left(\hat{x}^{(1)}_{i}\right)-V\left(x^{(1)}_{i}\right)+V\left(x^{(1)}_{i}\right)-\left[r\left(x^{(1)}_{i}\right)+\gamma V\left(x^{(2)}_{i}\right)\right]\right]^{2}\right]$
$=$ $~{}\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)-V\left(x^{(1)}_{i}\right)\right)^{2}\right]+$ (9)
$~{}2\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)-V\left(x^{(1)}_{i}\right)\right)\left(V\left(x^{(1)}_{i}\right)-\left[r\left(x^{(1)}_{i}\right)+\gamma V\left(x^{(2)}_{i}\right)\right]\right)\right]+$ (10)
$~{}\mathbb{E}\left[\left(V\left(x^{(1)}_{i}\right)-\left[r\left(x^{(1)}_{i}\right)+\gamma V\left(x^{(2)}_{i}\right)\right]\right)^{2}\right]$ (11)
The first summand (Equation 7) is equal to the IterVAML loss. Note that by our assumptions, $\hat{p}=p$ would minimize this perfectly.
To simplify the notation slightly, denote $\left[r\left(x^{(1)}_{i}\right)+\gamma V\left(x^{(2)}_{i}\right)\right]=V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)$. At the perfect model $p=\hat{p}$ the inner term of the expansion (Equation 7) is equal to 00, again because we assume that the value function is correct:
$\mathbb{E}\left[\underbrace{\left[V\left(\hat{x}^{(1)}_{i}\right)-V\left(x^{(1)}_{i}\right)\right]}_{=0}\left[V\left(x^{(1)}_{i}\right)-V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\right]\right].$
We now show that the final summand is not 0 and it is not independent of $\hat{p}$.
$\mathbb{E}\left[\left[V\left(x^{(1)}_{i}\right)-\left[\left[r\left(x^{(1)}_{i}\right)+\gamma V\left(x^{(2)}_{i}\right)\right]\right]\right]^{2}\right]$

### v3

@@ Appendix A Proofs and mathematical clarifications
The first proposition relies on the existence of a deterministic mapping, which we prove here as a lemma. The second proposition requires a statement over the minimizers of an equation with a variance-like term, we prove a general result as a lemma.
The proof of the first lemma relies heavily on several propositions from Bertsekas & Shreve (1978), which are restated in Subsection A.3 for reader’s convenience. Other topological statements are standard and can be find in textbooks such as Munkres (2018).
@@
As $\mathcal{Z}_{\mathcal{X}}$ is compact, the continuous function $V$ takes a maximum and minimum over the set $\mathcal{Z}_{\mathcal{X}}$. This follows from the compactness of $\mathcal{Z}_{\mathcal{X}}$ and the extreme value theorem. Furthermore $V_{\min}\leq\theta_{V,\mathcal{Z}}(z)\leq V_{\max}$ for every $z\in\mathcal{Z}_{\mathcal{X}}$. By the intermediate value theorem over compact, connected spaces, and the continuity of $V$, for every value $V_{\min}\leq v\leq V_{\max}$, there exists a $z\in\mathcal{Z}_{\mathcal{X}}$ so that $V(z)=v$.
Let $h:\mathcal{Z}_{\mathcal{X}}\times\mathcal{Z}_{\mathcal{X}}\rightarrow\mathbb{R}$ be the function $h(z,z^{\prime})=\left|\theta_{V,\mathcal{Z}}(z)-V(z^{\prime})\right|^{2}$. As $h$ is a composition of continuous functions, it is itself continuous. Let $h^{*}(z)=\min_{z^{\prime}\in\mathcal{Z}_{\mathcal{X}}}h(z,z^{\prime})$. For any $z\in\mathcal{Z}_{\mathcal{X}}$, by the intermediate value argument, there exist $z^{\prime}$ such that $V(z^{\prime})=v$. Therefore $h^{*}(z)$ can be minimized perfectly for all $z\in\mathcal{Z}_{\mathcal{X}}$.
Since $\mathcal{Z}_{\mathcal{X}}$ is compact, $h$ is defined over a compact subset of $\mathcal{Z}$. By Proposition 4, there exists a measurable function $f^{*}(z)$ so that $\min_{z^{\prime}}h(z,z^{\prime})=h(z,f^{*}(z))=0$. Therefore, the function $f^{*}$ has the property that $V(f^{*}(z))=\mathbb{E}_{p}\left[V(z^{\prime})|z\right]$, as this minimizes the function $h$.
@@
The following lemma shows that a function that minimizes a quadratic and a variance term cannot be the minimum function of the quadratic. This is used to show that the minimum of the MuZero value function learning term is not the same as applying the model-based Bellman operator.
Lemma 2
Let $g:\mathcal{X}\rightarrow\mathbb{R}$ be a function that is not constant almost everywhere and let $\mu$ be a non-degenerate probability distribution over $\mathcal{X}$. Let $\mathcal{F}$ be an open function space with $g\in\mathcal{F}$. Let $\mathcal{L}\left(f\right)=\mathbb{E}_{x\sim\mu}\left[\left(f(x)-g(x)\right)^{2}\right]+\mathbb{E}_{x\sim\mu}\left[f(x)g(x)\right]-\mathbb{E}_{x\sim\mu}\left[f(x)\right]\mathbb{E}_{\mu}\left[g(x)\right]$. Then $g\notin\operatorname*{arg\,min}_{f\in\mathcal{F}}\mathcal{L}(f)$.
Proof:
The proof follows by showing that there is a descent direction from $g$ that improves upon $\mathcal{L}$. For this, we construct the auxiliary function $\hat{g}(x)=g(x)-\epsilon g(x)$. Substituting $\hat{g}$ into $\mathcal{L}$ yields
$~{}\epsilon^{2}\mathbb{E}_{\mu}\left[g(x)^{2}\right]+\mathbb{E}_{\mu}\left[\left(g(x)-\epsilon g(x)\right)g(x)\right]-\mathbb{E}_{\mu}\left[\left(g(x)-\epsilon g(x)\right)\right]\mathbb{E}_{\mu}\left[g(x)\right]$
$=$ $~{}\epsilon^{2}\mathbb{E}_{\mu}\left[g(x)^{2}\right]+\left(1-\epsilon\right)\mathbb{E}_{\mu}\left[g(x)^{2}\right]-\left(1-\epsilon\right)\mathbb{E}_{\mu}\left[g(x)\right]^{2}.$
Taking the derivative of this function wrt to $\epsilon$ yields
$~{}\frac{\mathrm{d}}{\mathrm{d}\epsilon}\epsilon^{2}\mathbb{E}_{\mu}\left[g(x)^{2}\right]+\left(1-\epsilon\right)\mathbb{E}_{\mu}\left[g(x)^{2}\right]-\left(1-\epsilon\right)\mathbb{E}_{\mu}\left[g(x)\right]^{2}$
$=$ $~{}2\epsilon~{}\mathbb{E}_{\mu}\left[g(x)^{2}\right]-\mathbb{E}_{\mu}\left[g(x)^{2}\right]+\mathbb{E}_{\mu}\left[g(x)\right]^{2}.$
Setting $\epsilon$ to 0, obtain
$~{}\mathbb{E}_{\mu}\left[g(x)\right]^{2}-\mathbb{E}_{\mu}\left[g(x)^{2}\right]=\text{Var}_{\mu}\left[g(x)\right]$
By the Cauchy-Schwartz inequality, the variance is only 00 for a $g(x)$ constant almost everywhere. However, this violates the assumption. Therefore for any $\epsilon>0$ $\mathcal{L}\left(\hat{g}\right)\leq\mathcal{L}\left(g\right)$, due to the descent direction given by $-g(x)$. As we assume that the function class is open, there also exists an $\epsilon>0$ for which $g(x)-\epsilon g(x)\in\mathcal{F}$.
$\Box$
@@ A.1 Main propositions
Furthermore, the same $f^{*}$ is a minimizer of the population IterVAML loss:
@@
The rest of the proof follows standard approaches in textbooks such as Györfi et al. (2002). First, expand the equation to obtain:
@@
After expanding the square, we obtain three terms:
$\mathbb{E}\left[\left|{V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]}\right|^{2}\right]$
$+$ $2\mathbb{E}\left[\left[V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]\right]\left[\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]-V_{\mathcal{X}}(x^{\prime})\right]\right]$
$+$ $\mathbb{E}\left[\left|{\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]-V_{\mathcal{X}}(x^{\prime})}\right|^{2}\right]$
Apply the tower property to the inner term to obtain:
@@
$=$ $2\mathbb{E}\left[\left[V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]\right]\underbrace{\mathbb{E}\left[\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]-V_{\mathcal{X}}(x^{\prime})\middle|x^{\prime}\right]}_{=0}\right]=0$
Since the statement we are proving only applies to the minimum of the IterVAML loss, we will work with the $\operatorname*{arg\,min}$ of the loss function above. The resulting equation contains a term dependent on $f$ and one independent of $f$:
$\operatorname*{arg\,min}_{f}~{}$ $\mathbb{E}\left[\left|{V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]}\right|^{2}\right]+\mathbb{E}\left[\left|{\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]-V_{\mathcal{X}}(x^{\prime})}\right|^{2}\right]$
$=~{}\operatorname*{arg\,min}_{f}~{}$ $\mathbb{E}\left[\left|{V\left(f\left(\phi(x)\right)\right)-\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]}\right|^{2}\right].$
Finally, it is easy to notice that $V\left(f^{*}\left(\phi(x)\right)\right)=\mathbb{E}\left[V_{\mathcal{X}}(x^{\prime})\middle|x\right]$ by the definition of $f^{*}$. Therefore $f^{*}$ minimizes the final loss term and, due to that, the IterVAML loss.
@@
Assume a non-deterministic MDP with a fixed, but arbitrary policy $\pi$, and let $p$ be the transition kernel. Let $\mathcal{V}$ be an open set of functions, and assume that it is Bellman complete: $\forall V\in\mathcal{V}:\mathcal{T}V\in\mathcal{V}$.
@@
Notation:
For clarity of presentation denote samples from the real environment as $x^{(n)}$ for the $n$-th sample after a starting point $x^{(0)}$. This means that $x^{(n+1)}$ is drawn from $p\left(\cdot\middle|x^{(n)}\right)$. Similarly, $\hat{x}^{(n)}$ is the $n$-th sample drawn from the model, with $\hat{x}^{(0)}=x^{(0)}$. All expectations are taken over $x^{(0)}_{i}\sim\mu$ where $\mu$ is the data distribution, $\hat{x}^{(1)}_{i}\sim\hat{p}\left(\cdot\middle|x^{(0)}_{i}\right)$, $x^{(1)}_{i}\sim p\left(\cdot\middle|x^{(0)}_{i}\right)$, and $x^{(2)}_{i}\sim p\left(\cdot\middle|x^{(1)}_{i}\right)$. We use the tower property several times, all expectations are conditioned on ${x^{(0)}_{i}}_{i}$.
@@
By assumption, let $\hat{p}$ in the MuZero loss be the true transition kernel $p$. Expand the MuZero loss by $\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]$ and take its expectation:
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(\hat{p},\hat{V};\mathcal{D},V^{\prime})\right]$
@@
$~{}2~{}\mathbb{E}\left[\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)-\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)\right)\left(\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right)\right]+$ (5)
@@
We aim to study the minimizer of this term. The first term (Equation 4) is the regular bootstrapped Bellman residual with a target $V^{\prime}$. The third term (Equation 6) is independent of $\hat{V}$, so we can drop it when analyzing the minimization problem.
@@
as the remainder is independent of $\hat{V}$ again.
This remaining term however is not independent of $\hat{V}$ and not equal to 00 either. Instead, it decomposes into a variance-like term, using the conditional independence of $\hat{x}^{(1)}_{i}$ and $x^{(1)}_{i}$ given $x^{(0)}_{i}$:
@@
$=$ $~{}\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)\right]-\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right]$
$=$ $~{}\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)\right]-\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\right]\mathbb{E}\left[\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right].$
@@
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(p,\hat{V};\mathcal{D},V^{\prime})\right]$
$=$ $~{}\mathbb{E}\left[\left(\hat{V}\left(\hat{x}^{(1)}_{i}\right)-\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)\right)^{2}\right]+$
$~{}\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\left(\mathcal{T}V^{\prime}\right)\left(\hat{x}^{(1)}_{i}\right)\right]-\mathbb{E}\left[\hat{V}\left(\hat{x}^{(1)}_{i}\right)\right]\mathbb{E}\left[\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right].$
The first summand is the Bellman residual, which is minimized by $\mathcal{T}V^{\prime}$, which is in the function class by assumption. However, by Lemma 2, $\mathcal{T}{V^{\prime}}$ does not minimize the whole loss term under the conditions (open function class, non-constant value functions, and non-degenerate transition kernel).
$\Box$
Discussion:
The proof uses Bellman completeness, which is generally a strong assumption. However, this is only used to simplify showing the contradiction at the end, removing it does not remove the problems with the loss. The proof of Lemma 2 can be adapted to the case where $f(x)$ minimizes the difference to $g(x)$, instead of using $g(x)$ as the global minimum, but some further technical assumptions about the existence of minimizers and boundary conditions are needed. The purpose here is to show that even with very favorable assumptions such as Bellman completeness, the MuZero value function learning algorithm will not converge to an expected solution.
Similarly, the condition of openness of the function class simply ensures that there exists a function "nearby" that minimizes the loss better. This is mostly to remove edge cases, such as the case where the function class exactly contains the correct solution. Such cases, while mathematically valid, are uninteresting from the perspective of learning functions with flexible function approximations.
We only show the proof for the single step version and remove action dependence to remove notational clutter, the action-dependent and multi-step versions follow naturally.
A.2 Comparison of IterVAML and MuZero for model learning
If MuZero is instead used to only update the model $\hat{p}$, we obtain a similarity between MuZero and IterVAML. This result is similar to the one presented in Grimm et al. (2021), so we only present it for completeness sake, and not claim it as a fully novel contribution of our paper. While Grimm et al. (2021) show that the whole MuZero loss is an upper bound on an IterVAML-like term, we highlight the exact term the model learning component minimizes. However, we think it is still a useful derivation as it highlights some of the intuitive similarities and differences between IterVAML and MuZero and shows that they exist as algorithms on a spectrum spanned by different estimates of the target value function.
@@ A.2 Comparison of IterVAML and MuZero for model learning
We will choose a slightly different expansion than before, using $\mathbb{E}_{x_{i}^{(1)},x_{i}^{(2)}\sim p}\left[\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right]=\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]$
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(\hat{p},V;\mathcal{D},V^{\prime})\right]$
$=$ $~{}\mathbb{E}\left[\left[V\left(\hat{x}^{(1)}_{i}\right)-\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]+\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right]^{2}\right]$
$=$ $~{}\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)-\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]\right)^{2}\right]+$ (7)
$~{}2\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)-\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]\right)\left(\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right)\right]+$ (8)
[…]
The first summand (Equation 8) is similar to the IterVAML loss, instead of using the next state’s value function, the one-step bootstrap estimate of the Bellman operator is used.
The third term (Equation 8) is independent of $\hat{p}$ and can therefore be dropped. The second term decomposes into two terms again,
$~{}\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)-\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]\right)\left(\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right)\right]$
$=$ $~{}\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)\right)\left(\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right)\right]-$
[…]
The first summand is equal to 00, due to the conditional independence of $\hat{x}^{(1)}_{i}$ and $x^{(1)}_{i}$,
$~{}\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)\right)\left(\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]-\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right)\right]$
$=$ $~{}\mathbb{E}\left[V\left(\hat{x}^{(1)}_{i}\right)\right]\underbrace{\left(\mathbb{E}\left[\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]\right]-\mathbb{E}\left[\left[r\left(x^{(1)}_{i}\right)+\gamma V^{\prime}\left(x^{(2)}_{i}\right)\right]\right]\right)}_{=0}=0.$
The second remaining summand is independent of $\hat{p}$ and therefore irrelevant to the minimization problem.
Therefore, the MuZero model learning loss minimizes
$\mathbb{E}\left[\hat{\mathcal{L}}^{1}_{\text{MuZero}}(\hat{p},V;\mathcal{D},V^{\prime})\right]=\mathbb{E}\left[\left(V\left(\hat{x}^{(1)}_{i}\right)-\mathbb{E}\left[\mathcal{T}V^{\prime}\left(x^{(1)}_{i}\right)\right]\right)^{2}\right].$
In conclusion, the MuZero loss optimizes a closely related function to the IterVAML loss when used solely to update the model. There are three differences: First, the bootstrap value function estimator is used instead of the value function as the target value. Second, the current value function estimate is used for the model sample and the target network (if used) is applied for the bootstrap estimate. If the target network is equal to the value function estimate, this difference disappears. Finally, the loss does not contain the inner expectation around the model value function. This can easily be added to the loss and its omission in MuZero is unsurprising, as the loss was designed for deterministic environments and models.
The similarity between the losses suggests a potential family of decision-aware algorithms with different bias-variance characteristics, of which MuZero and IterVAML can be seen as two instances. It is also interesting to note that even without updating the value function, the MuZero loss performs an implicit minimization of the difference between the current value estimate and the Bellman operator via the model prediction. This is an avenue for further research, as it might explain some of the empirical success of the method disentangled from the value function update.
A.3 Propositions from Bertsekas & Shreve (1978)
For convenience, we quote some results from Bertsekas & Shreve (1978). These are used in the proof of Lemma 1.
Proposition 3 (Proposition 7.30 of Bertsekas & Shreve 1978)
Let $\mathcal{X}$ and $\mathcal{Y}$ be separable metrizable spaces and let $q(\mathrm{d}y|x)$ be a continuous stochastic kernel on $\mathcal{Y}$ given $\mathcal{X}$. If $f\in\mathcal{C}(\mathcal{X}\times\mathcal{Y})$, the function $\lambda:\mathcal{X}\rightarrow\mathbb{R}$ defined by
$\lambda(x)=\int f(x,y)q(\mathrm{d}y|x)$
is continuous.
Proposition 4 (Proposition 7.33 of Bertsekas & Shreve 1978)
[…]
$f^{*}(x)=\min_{y\in\mathcal{D}_{x}}f(x,y).$